# Unsupervised cell segmentation by fast Gaussian Processes

### v1

@@ 3.2 Binary cell masks
Figure 2: Comparison of binary image results across thresholds based on the absolute second difference in foreground pixels. The threshold, which ranges from 0-1, refers to the proportion of the maximum value of the predictive mean that is set as the binary cutoff. Images A, B, C, D, and E are the binary images generated by setting the corresponding threshold on cropped predictive mean image F. Note that the Image B corresponds with $\operatorname*{argmax}_{m}\Delta c_{k}(\alpha_{m})$ in Equation (12).
@@
Suppose each cropped image is $n_{k,1}\times n_{k,2}$ with predictive mean $\mathbf{F}^{*}_{k}$ with the $(i,j)$th entry being $F^{*}_{k}(\mathbf{x}_{i,j})$ and $\mbox{Vec}({\mathbf{F}^{*}_{k}})=\mathbf{f}^{*}_{k,v}$ given in Equation (11). Let $\{\alpha_{m}\}^{M}_{m=1}$ be a sequence of equally spaced threshold values from 0 to 1 with the default value $M=100$. For each threshold $\alpha_{m}$, we first compute the number of pixels $c_{k}(\alpha_{m})$ for which the normalized predictive mean exceeds each threshold value:
@@
where $\max(\mathbf{F}^{*}_{k})$ denotes the maximum value of the predictive mean $\mathbf{F}^{*}_{k}$ of the $k$th cropped image for $k=1,...,\tilde{K}$. Subsequently, we calculate the absolute difference $\Delta c_{k}(\alpha_{m})$ in pixel counts below
$\Delta c_{k}(\alpha_{m})=|c_{k}(\alpha_{m})-c_{k}(\alpha_{m-1})|.$
As $\Delta c_{k}(\alpha_{m})$ can contain noise, we compute the predictive mean of $\Delta c_{k}(\alpha_{m})$, denoted by $\Delta c^{*}_{k}(\alpha_{m})$, via a GP model with the default Matérn kernel in Equation (3) by the $\sf{RobustGaSP}$ package gu2018robustgasp .
We define $\alpha^{*}$ as the smallest threshold $\alpha_{m}$ after the point of maximum fluctuation, denoted as $\arg\max_{m}\Delta c_{k}(\alpha_{m})$, where the smoothed differences ${\Delta}c^{*}_{k}(\alpha_{m})$ stabilize within a specified tolerance relative to the variability of $\Delta c_{k}(\alpha_{m})$, an example of which is plotted in Figure 2. Formally, let $\tau_{k}$ represent the standard deviation of the differences ${\Delta}c^{*}_{k}(\alpha_{m})$. The optimal threshold $\alpha^{*}_{k}$ is given by:
$\alpha^{*}_{k}=\min_{m}\left\{\alpha_{m}:\left|{\Delta}c^{*}_{k}(\alpha_{m})-{\Delta}c^{*}_{k}(\alpha_{m-1})\right|<0.05\tau_{k},\,m>\arg\max_{m}\Delta c_{k}(\alpha_{m})\right\}.$ (12)
Figure 2 illustrates this process for determining the optimal threshold for segmenting cell objects from the background. While the first difference $\Delta c_{k}(\alpha_{m})$ measures the immediate change in the number of pixels above each threshold, the second difference more precisely captures the stabilization process of these changes. Initially, at extremely small thresholds, almost all pixels are classified as foreground or cell objects, leading to minimal changes in pixel counts between threshold values, shown as A in Figure 2. However, once the threshold starts to intersect the main background distribution, even small increments can trigger substantial variations, resulting in high second differences. As the threshold increases, the second difference reaches its maximum value at B in Figure 2 and then becomes erratic between B and C in Figure 2. This erratic behavior occurs because, in this transitional range, the predictive mean values near the decision boundary are very similar, so that even minor changes in the threshold result in disproportionately large shifts in the number of pixels classified as foreground or cell objects. Beyond this range, the second difference begins to stabilize and falls below a predetermined tolerance level show as E in Figure 2, indicating that further increases yield only negligible changes in segmentation. Any further increase in the threshold results in over-segmentation.
@@
for $k=1,...,\tilde{K}$. We complete panel (C) in Figure 1 by generating binary cell masks for all sub-images. To address a rare scenario with a higher-than-expected object count, which only happened once for all subimages, we implemented a re-thresholding step discussed in Supplementary Section S1. Finally, each object group is separated and detected via flood fill bhawnesh2020flood , as described in Section S2 in the Supplementary Material.
3.3 Labeled cell masks by the watershed algorithm
Figure 4: An image of two cell nuclei (upper row) and heights of negative distances to the nearest background pixels (lower row). (A) Heights of negative distances along the red straight line in the upper panel are plotted in the lower panel before the watershed algorithm starts. (B) At water level $=-13$, both catch basins are partially filled, as both local minima are less than the current water level. The two separate water sources have unique labels and are not yet touching. (C) At water level at around $-10.2$, the water sources from the two catch basins flow into each other and the watershed line is formed at water level at around $-10.2$. (D) At water level $=0$, the cell objects are filled with water, and the watershed operation is complete. Each cell is labeled and separated.
After obtaining the binary image, we apply the watershed algorithm to separate cell objects from $\sf{EBImage}$ package in $\sf{R}$ pau2010ebimage , which is particularly useful for segmenting objects connecting to each other. The watershed algorithm creates a topological representation of an image based on a distance map, which is derived from the binary image representing foreground objects (cells). The distance map assigns a value to each foreground pixel based on its Euclidean distance to the nearest background pixel; higher values represent larger distances from the background. These distance values act as heights in a topological landscape, providing a basis for the watershed function to simulate water through. The negative distance map is then calculated by taking the negative of the distance map. An example of the binary image and negative distance map is provided in Figure S2 in the Supplementary Material.
@@ 3.3 Labeled cell masks by the watershed algorithm
Figure 4 (B) shows the water levels at $-13$. Each catch basin is partially filled, as the water level is higher than the two local minima. As the water in each valley is still well separated, all emerged pixels in each valley are either labeled corresponding with $d_{1}$ or $d_{2}$, with no neighboring pixels containing a different label.
At water level around $-10.2$, shown in Figure 4 (C), the water from the two different basins begins to touch. The water from the two catch basins corresponding to $d_{1}$ and $d_{2}$ meet at the unlabeled pixel $p_{0}$ with a position close to 14, plotted in the x-coordinate. The Euclidean distances between $p_{0}$ and the local minima are utilized to determine its assignment, and a watershed line is formed. The Euclidean distance between $p_{0}$ and $d_{2}$ ($\approx 5.01$ units) is smaller than the distance between $p_{0}$ and $d_{1}$ ($\approx 5.62$ units), so $p_{0}$ is grouped with cell 2. The watershed line will serve as the boundary between the two cells.
At the water level 00, the objects are completely segmented, shown in Figure 4 (D). The watershed operation then halts, as this height is used as a stopping point since all cell objects have negative heights in the negative distance map. Additional details of the watershed algorithm, including definitions and edge cases, can be found in Vincent1991 .
@@
4 Simulated studies
In this section, we numerically compare our fast GPs with several alternative methods by simulated experiments. The comparison includes the fast GP of a Matérn kernel with roughness parameter being 2.5 in Equation (3) and the exponential kernels $K_{l}(d_{l})=\exp(-d_{l}/\gamma_{l})$ where $d_{l}$ is the distance of the $l$th coordinate of the input and $\gamma_{l}$ is a range parameter, for $l=1,2$. We also include three other alternative methods, the principal component analysis (PCA) berkooz1993proper , fast algorithm of multivariate Ornstein-Uhlenbeck processes (FMOU) lin2025fast , and dynamic mode decomposition (DMD) schmid2010dynamic . PCA is commonly used in image denoising by projecting image data onto low-dimensional spaces spanned by the leading eigenvectors of the data covariance matrix. FMOU provides a fast expectation-maximization (E-M) algorithm for parameter estimation in a latent factor model $\mathbf{Y}=\mathbf{U}_{0}\mathbf{Z}+\bm{\epsilon}$, where $\mathbf{U}_{0}$ is an orthogonal factor loading matrix and $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{d}]^{T}$ consists of $d$ independent latent processes, each modeled as an Ornstein-Uhlenbeck process with distinct correlation and variance parameters, and $\bm{\epsilon}$ are independent Gaussian noises. In each E-M iteration, parameters are updated using closed-form expressions. DMD is popular in video processing to extract the spatiotemporal structure in nonlinear dynamical systems of high-dimensions. It assumes a linear transition matrix $\mathbf{A}$, estimated as $\hat{\mathbf{A}}=\operatorname{argmin}_{\mathbf{A}}||\mathbf{Y}_{2:n_{2}}-\mathbf{A}\mathbf{Y}_{1:n_{2}-1}||$, where $\mathbf{Y}_{2:n_{2}}$ and $\mathbf{Y}_{1:n_{2}-1}$ represent the last and first $n_{2}-1$ columns of $\mathbf{Y}$, respectively. An efficient algorithm, known as exact DMD tu2014dmd , is proposed to identify the leading eigenpairs of $\mathbf{A}$ with computational complexity of $\min(\mathcal{O}(n_{1}n_{2}^{2}),\mathcal{O}(n_{1}^{2}n_{2}))$.
@@ 4 Simulated studies
To compare the model performance in noising filtering, we consider the error of the predictive mean of the noisy observations, quantified by the root of mean squared error (RMSE):
@@
Figure 8: Comparison of the computational time in seconds for evaluating the profile likelihood of $\mathbf{Y}_{v}$ using direct and fast computation under varying sample sizes $N$. The direct computation with the exponential kernel (Direct-Exp) is plotted as orange dots; the direct computation with the Matérn kernels in Equation (3) (Direct-Mat) is plotted as green triangles; 3) the fast computation with the exponential kernel (Fast-Exp) is plotted as pink squares; The fast computation with the Matérn kernels in Equation (3) (Fast-Mat) is plotted as blue crosses. The left panel shows the computational time in seconds (original scale), while the right panel displays the logarithmic scale.
@@
5 Real data analysis
@@ 5.1 Criteria and methods
Figure 9: Results for segmenting cell nuclei. (A) AP scores for GP-based segmentation and ImageJ across different thresholds. (B) Boxplots of AP scores for each method. (C) True boundaries of the fifth test image with 1024 $\times$ 1024 pixels. (D) Boundaries generated by the GP-based method for the fifth test image. (E) Boundaries generated using ImageJ for the fifth test image.
@@ 6 Concluding remarks
In this study, we have presented a scalable cell image segmentation method for both the cell nucleus and cytoplasm. By leveraging fast Gaussian processes, automated thresholding and watershed operations, we shown the process is not only more automated, but the results are more precise than other unsupervised segmentation methods, such as ImageJ. A key strength of the proposed segmentation algorithm is that it does not rely on parameter tuning or labelled training datasets. The results of our unsupervised method demonstrate versatility across different cell channels, which makes it an appealing option for more general segmentation tasks beyond cell segmentation.
Several future directions are worth exploring. First, modern microscopes provide time-lapse videos for tracking the changes in cellular behavior over time. A future goal is to integrate our segmentation methods with particle tracking and linking algorithms that solve a linear assignment problem jaqaman2008robust . As our approach does not restrict the shape of the object, it can be extended for image segmentation with objects undergoing dynamical changes, including microscopy images for the nucleation and growth process in crystallization wu2022nucleation and protein dynamics lee2021myosin . Second, when labelled image data are available, it is worthwhile to explore how these data can be used for constraining the shapes of the objects to improve estimation and simultaneously enable segmenting new objects not in the database. Third, the majority of cell segmentation methods are developed for 2D microscopy images. It is of interest to develop cell segmentation methods for microscopy that capture cell behaviors in a 3D environment, such as the orientational order and alignment huang2025cell . Furthermore, the uncertainty is often not quantified by the image segmentation methods, as the segmentation process often contains multiple steps, and the uncertainty from the convolutional neural network is an open question. As we model the image data with a probabilistic model, a future goal is to propagate the uncertainty throughout the analysis for uncertainty quantification.
@@ Acknowledgement
MG acknowledges partial support from the National Science Foundation under Award No. OAC-2411043. HY acknowledges partial support from NSF under the Materials Research Science and Engineering Center (MRSEC) at UC Santa Barbara under Award No. DMR-2308708.
@@
$\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{y}_{v}$ $=\mathbf{1}^{T}_{N}(\mathbf{Q}\tilde{\bm{\Lambda}}\mathbf{Q}^{T})^{-1}\mathbf{y}_{v}$
@@
$\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{1}_{N}$ $=\mathbf{1}^{T}_{N}(\mathbf{U}_{2}\otimes\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}\mathbf{(}\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\mathbf{1}_{N}$
@@
$=[(\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\mbox{Vec}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})]^{T}\tilde{\bm{\Lambda}}^{-1}(\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\mbox{Vec}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})$
$=[\mbox{Vec}(\mathbf{U}_{1}^{T}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})\mathbf{U}_{2})]^{T}\tilde{\bm{\Lambda}}^{-1}\mbox{Vec}(\mathbf{U}_{1}^{T}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})\mathbf{U}_{2})$
@@
$=\hat{\mu}\mathbf{1}_{N}+(\mathbf{U}_{2}\otimes\mathbf{U}_{1})(\bm{\Lambda}_{2}\otimes\bm{\Lambda}_{1})(\mathbf{U}^{T}_{2}\otimes\mathbf{U}^{T}_{1})(\mathbf{U}_{2}\otimes\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}\mathbf{(}\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})$
$=\hat{\mu}\mathbf{1}_{N}+(\mathbf{U}_{2}\otimes\mathbf{U}_{1})[(\bm{\Lambda}_{2}\otimes\bm{\Lambda}_{1})\tilde{\bm{\Lambda}}^{-1}]\mbox{Vec}(\mathbf{U}_{1}^{T}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})\mathbf{U}_{2})$
@@
$\mathbf{R}\mathbf{\tilde{R}}^{-1}\mathbf{R}$ $=(\mathbf{R}_{2}\otimes\mathbf{R}_{1})(\mathbf{U}_{2}\otimes\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}\mathbf{(}\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})(\mathbf{R}_{2}\otimes\mathbf{R}_{1})$
@@
$\hat{\sigma}^{2}c^{*}_{i,j}$ $=\hat{\sigma}^{2}\left(1-(\mathbf{r}^{T}_{j,2}\mathbf{U}_{2}\otimes\mathbf{r}^{T}_{i,1}\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}(\mathbf{U}^{T}_{2}\mathbf{r}_{j,2}\otimes\mathbf{U}^{T}_{1}\mathbf{r}_{i,1})\right)$

### v2

@@ 3.2 Generating binary cell masks
Figure 2: Comparison of binary image results across thresholds based on the absolute second difference in foreground pixels. The threshold, which ranges from 0-1, refers to the proportion of the maximum value of the predictive mean that is set as the binary cutoff. Images A, B, C, D, and E are the binary images generated by setting the corresponding threshold on cropped predictive mean image F. Note that the Image B corresponds with $\operatorname*{argmax}_{m}\Delta c_{k}(\alpha_{m})$ in Equation (3.2).
@@
Suppose each cropped image is $n_{k,1}\times n_{k,2}$ with predictive mean $\mathbf{F}^{*}_{k}$ with the $(i,j)$th entry being $F^{*}_{k}(\mathbf{x}_{i,j})$ and $\mbox{Vec}({\mathbf{F}^{*}_{k}})=\mathbf{f}^{*}_{k,v}$ given in Equation (11). As the intensity values of the pixels are between 0 to 1 and the predictive uncertainty is small, the range of $F^{*}_{k}(\mathbf{x}_{i,j})$ is also approximately between 0 to 1. As the image has only one type of cells, the change of pixel values $F^{*}_{k}(\mathbf{x}_{i,j})$ is large if $\mathbf{x}_{i,j}$ is the pixel at the boundary of a cell object, and small elsewhere. Thus we first compute the normalized predictive mean exceeds each threshold value:
@@
where $\{\alpha_{m}\}^{M}_{m=1}$ a sequence of equally spaced thresholds from 0 to 1 with the default value $M=100$, and $\max(\mathbf{F}^{*}_{k})$ denotes the maximum value of the predictive mean $\mathbf{F}^{*}_{k}$ of the $k$th cropped image for $k=1,...,\tilde{K}$. In general, the range of $\{\alpha_{m}\}^{M}_{m=1}$ can be chosen to be the range of $F^{*}_{k}(\mathbf{x}_{i,j})/\max(\mathbf{F}^{*}_{k})$. Subsequently, we calculate the difference $\Delta c_{k}(\alpha_{m})$ in pixel counts below
$\Delta c_{k}(\alpha_{m})={c_{k}(\alpha_{m-1})-c_{k}(\alpha_{m})}.$
As $\Delta c_{k}(\alpha_{m})$ may not be smooth over $\alpha_{m}$, we compute the predictive mean of $\Delta c_{k}(\alpha_{m})$, denoted by $\Delta c^{*}_{k}(\alpha_{m})$, via a GP model with the default Matérn kernel in Equation (3) by the $\sf{RobustGaSP}$ package gu2018robustgasp .
We define $\alpha^{*}$ as the smallest threshold $\alpha_{m}$ after the point of maximum fluctuation, denoted as $\arg\max_{m}\Delta c^{*}_{k}(\alpha_{m})$, where the smoothed differences ${\Delta}c^{*}_{k}(\alpha_{m})$ stabilize within a specified tolerance relative to the variability of $\Delta c_{k}(\alpha_{m})$, an example of which is plotted in Figure 2. Formally, let $\tau_{k}$ represent the standard deviation of the differences ${\Delta}c^{*}_{k}(\alpha_{m})$. The optimal threshold $\alpha^{*}_{k}$ is given by:
$\alpha^{*}_{k}=$ $\min_{m}\{\alpha_{m}:\left|{\Delta}c^{*}_{k}(\alpha_{m})-{\Delta}c^{*}_{k}(\alpha_{m-1})\right|<0.05\tau_{k},\,$
$m>\arg\max_{m}\Delta c^{*}_{k}(\alpha_{m})\}.$ (12)
Figure 2 illustrates this process for determining the optimal threshold for segmenting cell objects from the background. While the first difference $\Delta c_{k}(\alpha_{m})$ measures the immediate change in the number of pixels above each threshold, the second difference more precisely captures the stabilization process of these changes. Initially, at extremely small thresholds, almost all pixels are classified as foreground or cell objects, leading to minimal changes in pixel counts between threshold values, shown as A in Figure 2. However, once the threshold starts to intersect the main background distribution, even small increments can trigger substantial variations, resulting in high second differences. As the threshold increases, the second difference reaches its maximum value at B in Figure 2 and then becomes erratic between B and C in Figure 2. This erratic behavior occurs because, in this transitional range, as the predictive mean values near the decision boundary are very similar, even minor changes in the threshold result in disproportionately large shifts in the number of pixels classified as foreground or cell objects. Beyond this range, the second difference begins to stabilize and falls below a predetermined tolerance level show as E in Figure 2, indicating that further increases yield only negligible changes in segmentation. Any further increase in the threshold results in over-segmentation.
@@
for $k=1,...,\tilde{K}$. We complete panel (C) in Figure 1 by generating binary cell masks for all sub-images. To address a rare scenario with a higher-than-expected object count, which only happened once for all sub-images, we implemented a re-thresholding step discussed in Supplementary Section S3. Finally, each object group is separated and detected via flood fill bhawnesh2020flood , as described in Section S4 in the Supplementary Material.
3.3 Labeling cell masks by the watershed algorithm
Figure 4: An image of two cell nuclei (upper row) and heights of negative distances to the nearest background pixels (lower row). (A) Heights of negative distances along the red straight line in the upper panel are plotted in the lower panel before the watershed algorithm starts. (B) At water level =13=-13, both catch basins are partially filled, as both local minima are less than the current water level. The two separate water sources have unique labels and are not yet touching. (C) At water level at around 10.2-10.2, the water sources from the two catch basins flow into each other and the watershed line is formed at water level at around 10.2-10.2. (D) At water level =0=0, the cell objects are filled with water, and the watershed operation is complete. Each cell is labeled and separated.
After obtaining the binary image, we apply the watershed algorithm to separate cell objects from $\sf{EBImage}$ package in $\sf{R}$ pau2010ebimage , which is particularly useful for segmenting objects connecting to each other. The watershed algorithm creates a topological representation of an image based on a distance map, which is derived from the binary image representing foreground objects (cells). The distance map assigns a value to each foreground pixel based on its Euclidean distance to the nearest background pixel; higher values represent larger distances from the background. These distance values act as heights in a topological landscape, providing a basis for the watershed function to simulate water through. The negative distance map is then calculated by taking the negative of the distance map. Examples of the binary image and negative distance map are provided in Section S5 in the Supplementary Material.
@@ 3.3 Labeling cell masks by the watershed algorithm
Figure 4 (B) shows the water levels at 13-13. Each catch basin is partially filled, as the water level is higher than the two local minima. As the water in each valley is still well separated, all emerged pixels in each valley are either labeled corresponding with $d_{1}$ or $d_{2}$, with no neighboring pixels containing a different label.
At water level around 10.2-10.2, shown in Figure 4 (C), the water from the two different basins begins to touch. The water from the two catch basins corresponding to $d_{1}$ and $d_{2}$ meet at the unlabeled pixel $p_{0}$ with a position close to 14, plotted in the x-coordinate. The Euclidean distances between $p_{0}$ and the local minima are utilized to determine its assignment, and a watershed line is formed. The Euclidean distance between $p_{0}$ and $d_{2}$ ($\approx 5.01$ units) is smaller than the distance between $p_{0}$ and $d_{1}$ ($\approx 5.62$ units), so $p_{0}$ is grouped with cell 2. The watershed line will serve as the boundary between the two cells.
At the water level 0, the objects are completely segmented, shown in Figure 4 (D). The watershed operation then halts, as this height is used as a stopping point since all cell objects have negative heights in the negative distance map. Additional details of the watershed algorithm, including definitions and edge cases, can be found in Vincent1991 .
@@
4 Numerical studies of image denoising
[…]
@@ 4 Numerical studies of image denoising
To compare the model performance in noising filtering, we consider the error of the predictive mean of the noisy observations, quantified by the root mean squared error (RMSE):
@@
Figure 8: Comparison of the computational time in seconds for evaluating the profile likelihood of $\mathbf{Y}_{v}$ using direct and fast computation under varying sample sizes NN. The direct computation with the exponential kernel (Direct-Exp) is plotted as orange dots; the direct computation with the Matérn kernels in Equation (3) (Direct-Mat) is plotted as green triangles; 3) the fast computation with the exponential kernel (Fast-Exp) is plotted as pink squares; The fast computation with the Matérn kernels in Equation (3) (Fast-Mat) is plotted as blue crosses. The left panel shows the computational time in seconds (original scale), while the right panel displays the logarithmic scale.
@@
5 Numerical studies of cell segmentation
@@ 5.1 Criteria and methods
Figure 9: Results for segmenting cell nuclei. (A) AP scores for GP-based segmentation and ImageJ across different thresholds. (B) Boxplots of AP scores for each method. These plots are created using the APs for each of the 5 images using each method across different thresholds. (C) True boundaries of the fifth test image with $1024\times 1024$ pixels. (D) Boundaries generated by the GP-based method for the fifth test image. (E) Boundaries generated using ImageJ for the fifth test image.
@@ 6 Concluding remarks
In this study, we have presented a scalable cell image segmentation method for both the cell nucleus and cytoplasm. By leveraging fast Gaussian processes, automated thresholding and watershed operations, we shown the process is not only more automated, but the results are more precise than other unsupervised segmentation methods, such as ImageJ. A key strength of the proposed segmentation algorithm is that it does not rely on parameter tuning or labeled training datasets. The results of our unsupervised method demonstrate versatility across different cell channels, making it an appealing option for more general segmentation tasks beyond cell segmentation.
Several future directions are worth exploring. First, modern microscopes provide time-lapse videos for tracking the changes in cellular behavior over time. A future goal is to integrate our segmentation methods with particle tracking and linking algorithms that solve a linear assignment problem jaqaman2008robust . As our approach does not restrict the shape of the object, it can be extended for image segmentation with objects undergoing dynamical changes, including microscopy images for the nucleation and growth process in crystallization wu2022nucleation and protein dynamics lee2021myosin . Second, when labeled image data are available, it is worthwhile to explore how these data can be used for constraining the shapes of the objects to improve estimation and simultaneously enable segmenting new objects not in the database. Third, the majority of cell segmentation methods are developed for 2D microscopy images. It is of interest to develop cell segmentation methods for microscopy that capture cell behaviors in a 3D environment, such as the orientational order and alignment huang2025cell . Similar to some popular cell segmentation tools, such as Cellpose stringer2021cellpose , our approach aims to segment images that contain the same type of cell. A future direction is to leverage techniques in tools such as CellProfiler carpenter2006cellprofiler to segment images that contain multiple types of cells. Furthermore, the uncertainty is often not quantified by the image segmentation methods, as the segmentation process often contains multiple steps, and the uncertainty from the convolutional neural network is an open question. As we model the image data with a probabilistic model, a future goal is to propagate the uncertainty throughout the analysis for uncertainty quantification.
@@ Acknowledgement
This research is partially supported by the Materials Research Science and Engineering Center (MRSEC) by the National Science Foundation under Award No. DMR-2308708 and the Cyberinfrastructure for Sustained Scientific Innovation program by the National Science Foundation under Award No. OAC-2411043. We thank three anonymous referees for their comments that substantially improved this article.
Supplementary Material
The supplementary material provides additional details for image segmentation, experiments, and generation of ground truth.
@@
$\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{y}_{v}$
$=$ $\mathbf{1}^{T}_{N}(\mathbf{Q}\tilde{\bm{\Lambda}}\mathbf{Q}^{T})^{-1}\mathbf{y}_{v}$
@@
$\mathbf{1}^{T}_{N}\mathbf{\tilde{R}}^{-1}\mathbf{1}_{N}$
$=$ $\mathbf{1}^{T}_{N}(\mathbf{U}_{2}\otimes\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}\mathbf{(}\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\mathbf{1}_{N}$
@@
$=$ $[(\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\mbox{Vec}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})]^{T}$
$\tilde{\bm{\Lambda}}^{-1}\times(\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\mbox{Vec}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})$
$=$ $[\mbox{Vec}(\mathbf{U}_{1}^{T}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})\mathbf{U}_{2})]^{T}$ (17)
$\tilde{\bm{\Lambda}}^{-1}\mbox{Vec}(\mathbf{U}_{1}^{T}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})\mathbf{U}_{2})$
@@
$=$ $\hat{\mu}\mathbf{1}_{N}+(\mathbf{U}_{2}\otimes\mathbf{U}_{1})(\bm{\Lambda}_{2}\otimes\bm{\Lambda}_{1})(\mathbf{U}^{T}_{2}\otimes\mathbf{U}^{T}_{1})\times$
$(\mathbf{U}_{2}\otimes\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}\mathbf{(}\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})(\mathbf{y}_{v}-\hat{\mu}\mathbf{1}_{N})$
$=$ $\hat{\mu}\mathbf{1}_{N}+(\mathbf{U}_{2}\otimes\mathbf{U}_{1})[(\bm{\Lambda}_{2}\otimes\bm{\Lambda}_{1})\tilde{\bm{\Lambda}}^{-1}]\times$
$\mbox{Vec}(\mathbf{U}_{1}^{T}(\mathbf{Y}-\hat{\mu}\mathbf{1}_{n_{1}\times n_{2}})\mathbf{U}_{2})$
@@
$\mathbf{R}\mathbf{\tilde{R}}^{-1}\mathbf{R}=$ $(\mathbf{R}_{2}\otimes\mathbf{R}_{1})(\mathbf{U}_{2}\otimes\mathbf{U}_{1})\tilde{\bm{\Lambda}}^{-1}\mathbf{(}\mathbf{U}_{2}^{T}\otimes\mathbf{U}_{1}^{T})\times$
$(\mathbf{R}_{2}\otimes\mathbf{R}_{1})$
@@
$\hat{\sigma}^{2}c^{*}_{i,j}=$ $\hat{\sigma}^{2}-\hat{\sigma}^{2}(\mathbf{r}^{T}_{j,2}\mathbf{U}_{2}\otimes\mathbf{r}^{T}_{i,1}\mathbf{U}_{1})\times$
$\tilde{\bm{\Lambda}}^{-1}(\mathbf{U}^{T}_{2}\mathbf{r}_{j,2}\otimes\mathbf{U}^{T}_{1}\mathbf{r}_{i,1})$